A 30-year maturity bond making annual coupon payments with a coupon rate of 7.5% has duration of 12.27 years and convexity of 216.28. The bond currently sells at a yield to maturity of 8%.

e-1. Find the price of the bond if its yield to maturity increases to 9%. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

e-2. What price would be predicted by the duration rule? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

e-3. What price would be predicted by the duration-with-convexity rule? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

e-4. What is the percent error for each rule? (Negative answers should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to 2 decimal places.
Answer #1

e-1

New bond price @ YTM =9
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =30
Bond Price =∑ [(7.5*1000/100)/(1 + 9/100)^k]     +   1000/(1 + 9/100)^30
                   k=1
Bond Price = 845.9

e-2

New bond price @ YTM =9 using duration
Mod.duration prediction = -Mod. Duration*Yield_Change*Bond_Price
=-12.27*0.01*943.71
=-115.793217
New bond price = bond price+Modified duration prediction
=943.71+-115.793217
=827.92

e-3

New bond price @ YTM =9 using duration and convexity
Convexity adjustment = 0.5*convexity*Yield_Change^2*Bond_Price
=0.5*216.28*0.01^2*943.71
=10.20527994
New bond price = bond price+Mod.duration pred.+convex. Adj.
=943.71+-115.79+10.21
=838.13

e-4

Percentage error for YTM =9 and duration rule
Difference in price predicted and actual
=predicted price-actual price
=827.916783-845.9
=-17.983
%age difference = difference/actual
=-17.983/845.9
=-2.13%
Percentage error for YTM =9 and duration & convexity rule
Difference in price predicted and actual
=predicted price-actual price
=838.13-845.9
=-7.77
%age difference = difference/actual
=-7.77/845.9
=-0.92%